### v1

@@ II Methodology
The particular bin under consideration, for which we have calculated the binned bispectrum $\hat{B}^{s}(k_{1\perp},k_{2\perp},k_{3\perp},k_{1\parallel},k_{2\parallel})$, considers all closed triangles $({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ that have $(k_{1\parallel},k_{2\parallel},k_{3\parallel})$ fixed, and $({\bm{k}}_{1\perp},{\bm{k}}_{2\perp},{\bm{k}}_{3\perp})$ lie within the annular rings $(a_{1},a_{2},a_{3})$ respectively. Note that the binned bispectrum is a function of 5 independent parameters $(k_{1\perp},k_{2\perp},k_{3\perp},k_{1\parallel},k_{2\parallel})$, that describes the mean shape, size, and orientation with respect to $\hat{z}$. We use Eq. 2 and Eq. 3 to transform from $(k_{1\perp},k_{2\perp},k_{3\perp},k_{1\parallel},k_{2\parallel})$ to $(\hat{\textbf{p}},k_{1},\mu,t)$, whereby we have the binned bispectrum estimator $\hat{B}^{s}(\hat{\bm{p}},k_{1},\mu,t)$ that we use in Eq. 5 to calculate the multipole moments $\hat{\bar{B}}_{\ell}^{m}(k_{1},\mu,t)$. The resulting values of $(k_{1},\mu,t)$ are not uniformly spaced, and we have further binned $\hat{\bar{B}}_{\ell}^{m}(k_{1},\mu,t)$ so as to uniformly sample the $(k_{1},\mu,t)$ parameter space. The bin averaged values of $\hat{\bar{B}}_{\ell}^{m}({k}_{1},{\mu},{t})$ are computed as the weighted mean using the number of triangles $N_{\rm tri}$ as weights.
@@ III.1 Theoretical expressions
To validate the estimator, we generate a mildly non-Gaussian random field in redshift space $\delta^{s}(\bm{x})$ for which the bispectrum multipole moments are analytically derived. We start with generating a Gaussian random field in real space $\delta^{r}_{\rm G}(\bm{x})$ using a specified power spectrum $P(k)$ and introduce non-Gassianity through a local quadratic bias,
@@
$B^{r}(k_{1},k_{2},k_{3})=2f_{\rm NG}[P(k_{1})P(k_{2})+P(k_{2})P(k_{3})+P(k_{3})P(k_{1})]~{}.$ (14)
The field is mapped to redshift space by applying linear RSD,
@@
where $\Delta^{r}({\bm{k}})$ and $\Delta^{s}({\bm{k}})$ are the Fourier transform of the field in real and redshift space respectively and $\beta_{1}$ is the linear RSD parameter. The linear RSD theory predicts the redshift space bispectrum $B^{s}(\hat{\textbf{p}},k_{1},\mu,t)$ for this field to be,
$B^{s}(\hat{\textbf{p}},k_{1},\mu,t)=(1+\beta_{1}\mu_{1}^{2})(1+\beta_{1}\mu_{2}^{2})(1+\beta_{1}\mu_{3}^{2})B^{r}(k_{1},k_{2},k_{3})~{}.$ (16)
The analytical expressions for the multipole moments $B_{\ell}^{m}(k_{1},\mu,t)$ can be calculated using Eq.(4). These expressions are rather lengthy and hence not explicitly shown here. We refer the reader to Eqs. (24-29,31-33,A1-A7) of [25] for details. The analytical expressions are expected to hold true for sufficiently small values of $f_{\rm NG}$ such that $f_{\rm NG}\,\sigma_{\rm G}\ll 1$, where ${\sigma^{2}_{\rm G}=\langle[\delta^{r}_{\rm G}]^{2}\rangle}$ is the variance of ${\delta^{r}_{\rm G}}(\bm{x})$.
@@ III.2 Simulations
We simulated the field in a box of volume $V=L^{3}=(215~{}{\rm Mpc})^{3}$ divided in $N_{\rm g}=384$ grids in each direction. The inter-grid spacing for the box comes out to be $0.56~{}{\rm Mpc}$ and $k_{\rm min}=0.029~{}{\rm Mpc}^{-1}$. We use the input power spectrum $P(k)=k^{-2}$ to generate $\delta_{\rm G}^{r}$. The non-Gaussinity parameter is chosen to be $f_{\rm NG}=0.5$ for which $f_{\rm NG}\sigma_{\rm G}\approx 0.2$ and set linear RSD parameter $\beta_{1}=1$. We have generated $50$ independent realizations of the field and used these to estimate the mean $\bar{B}_{\ell}^{m}(k_{1},\mu,t)=\langle\hat{\bar{B}}_{\ell}^{m}(k_{1},\mu,t)\rangle$ and also the $1\sigma$ error bars.
@@ III.3 Binning and discreteness effects
In this work, we have validated the estimator by comparing the values of $\bar{B}_{\ell}^{m}(k_{1},\mu,t)$ estimated from the simulations with the analytical predictions for the same quantities. However, the analytical predictions $[\bar{B}_{\ell}^{m}]_{\rm c}$ presented in [25], which have been obtained by using Eq.(16) in Eq.(4), assume the continuum limit (as denoted by the subscript ${\rm c}$). In reality, we only have discrete ${\bm{k}}$ modes, and the integral in Eq.(4) is replaced by a sum in Eq.(5). We further bin the estimated $\bar{B}_{\ell}^{m}(k_{1},\mu,t)$ in $(k_{1},\mu,t)$ to increase the SNR and also obtained values at regular intervals of $(k_{1},\mu,t)$. For this, we introduce the discrete analytical predictions $[\bar{B}_{\ell}^{m}]_{\rm d}$ (as denoted by the subscript ${\rm d}$), which has been calculated by using the analytical predictions Eq.(16) in Eq.(5). The resulting predictions have also been binned in exactly the same way as for the estimator. In this work, we have validated the estimator by comparing $\bar{B}_{\ell}^{m}(k_{1},\mu,t)$ with the analytical predictions $[\bar{B}_{\ell}^{m}]_{\rm d}$.
Most theories which model the bispectrum, typically provide analytical predictions for $[\bar{B}_{\ell}^{m}]_{\rm c}$. Therefore, it is often desirable to compare the estimated values $\bar{B}_{\ell}^{m}(k_{1},\mu,t)$ with the analytical predictions $[\bar{B}_{\ell}^{m}]_{\rm c}$ and not $[\bar{B}_{\ell}^{m}]_{\rm d}$. A few earlier works [31, 20] have addressed this issue by introducing ‘discreteness weights’ $w_{\ell}^{m}(k_{1},\mu,t)$
@@ III.4 Results
The $[\bar{B}_{\ell}^{m}]_{\rm c}$ values are predicted to have a scale-invariant power-law $k_{1}$ dependence, and it is thus adequate to show the results for just a single value of $k_{1}$. The number of triangles scales as $N_{\rm tri}\propto k_{1}^{6}$ [36], and we have chosen a large value $k_{1}=2.1~{}{\rm Mpc}^{-1}$ with a bin width $dk_{1}=0.2~{}{\rm Mpc}^{-1}$ so as to achieve a large value of $N_{\rm tri}$ and obtain a reliable (high signal to noise ratio: SNR) estimate of $\bar{B}_{\ell}^{m}$. So as to accurately capture the shape dependence of $\bar{B}_{\ell}^{m}(k_{1},\mu,t)$, we have divided the $(\mu,t)$ plane into fine bins of width $(d\mu,dt)=(0.04,0.04)$ for evaluating the $\ell=0$ and $2$ multipoles. The values of $\bar{B}_{\ell}^{m}$ and the SNR are smaller for the higher multipoles $\ell>2$, and we have used coarser $(\mu,t)$ bins for these.
@@
Fig.2 illustrates the $(\mu,t)$ bins that we have used to parameterize the triangle shape dependence of $\bar{B}_{\ell}^{m}(k_{1},\mu,t)$ for $\ell\leq 2$. It also shows $N_{\rm tri}$ the number of triangles corresponding to each bin. Considering the $\mu-t$ plane, the $\mu=1$ boundary corresponds to linear triangles where the two larger sides of the triangle nearly aligned in the same direction. The $t=1$ and $12\mu t=$ 1 borders correspond to the L and S Isosceles triangles respectively. The two largest sides are equal for L Isosceles triangles, whereas the two smaller sides are equal for S. We see that in all cases the number of triangles is of the order of $10^{8}$, and it is largest for L Isosceles triangles. Further, $N_{\rm tri}$ decreases almost monotonically with decline in $t$, i.e., $N_{\rm tri}\propto t$, and the number falls by a factor of $10$ as triangles are deformed to the stretched limit. Nonetheless, $N_{\rm tri}$ does not vary significantly with $\mu$.
@@
Figure 3: Results for $\ell=0$ and $2$ multipoles. It shows the estimates and their deviation from analytical prediction for all triangle shapes ($\mu,t$) having a fixed size $k_{1}=2.1~{}{\rm Mpc}^{-1}$. The uppermost row depicts the results for the monopole ($\ell=0,m=0$). The second, third, and fourth rows show $\ell=2$ and $m=0,1,2$ moments, respectively. The first column illustrates the mean estimates $\bar{B}_{\ell}^{m}$ from simulations. The second column shows $\Delta_{\ell}^{m}=|\Delta{B}_{\ell}^{m}|/{\sigma}{B_{\ell}^{m}}$, $\Delta{B}_{\ell}^{m}$ is the deviation of the estimates from analytical predictions and ${\sigma}B_{\ell}^{m}$ is the expected r.m.s. statistical fluctuation. The third column explicitly shows the estimated ${B}_{\ell}^{m}$ values (Red points with $5{\sigma}{B_{\ell}^{m}}$ error bars) for linear, L, and S Isosceles triangles. The various lines show the corresponding analytical predictions. Results are plotted against $\mu$ for L and S Isosceles triangles, and against $t$ for linear triangles.
@@
Fig.3 shows the results for the $\ell=0$ and $2$ multipoles. Here, each row corresponds to a different combination of $\ell$ and $m$, as indicated in the figure. The panels in the left column show $\bar{B}_{\ell}^{m}$ estimated from the simulations. The analytical predictions $[\bar{B}_{\ell}^{m}]_{\rm d}$ are visually indistinguishable from the $\bar{B}_{\ell}^{m}$ shown here, and we have not explicitly shown $[\bar{B}_{\ell}^{m}]_{\rm d}$. The panels in the middle column show $\Delta_{\ell}^{m}=\mid\Delta{B}_{\ell}^{m}\mid/\sigma{B_{\ell}^{m}}$ where $\Delta\bar{B}_{\ell}^{m}={\bar{B}}_{\ell}^{m}-[\bar{B}_{\ell}^{m}]_{\rm d}$ is the difference between the estimated value and the analytical predictions, and $\sigma{B_{\ell}^{m}}$ (estimated from the $50$ realizations of the simulations) is the r.m.s. statistical fluctuations expected in $\Delta\bar{B}_{\ell}^{m}$. $\Delta_{\ell}^{m}$ provides an estimate of whether the deviations between the estimator and analytical predictions are consistent with the expected statistical fluctuations or not. The panels in the right column show a comparison of $\bar{B}_{\ell}^{m}$ and $[\bar{B}_{\ell}^{m}]_{\rm d}$ along the three boundaries of the $\mu-t$ plane.

### v2

@@ II Methodology
The particular bin under consideration, for which we have calculated the binned bispectrum $\hat{B}^{s}(k_{\perp r_{1}},k_{\perp r_{2}},k_{\perp r_{3}},k_{\parallel n_{1}},k_{\parallel n_{2}})$ considers all closed triangles where the three sides, namely $({\bm{u}},k_{\parallel n_{1}})$, $({\bm{v}},k_{\parallel n_{2}})$ and $({\bm{w}},k_{\parallel n_{3}})$, have $(k_{\parallel n_{1}},k_{\parallel n_{2}},k_{\parallel n_{3}})$ fixed, whereas $({\bm{u}},{\bm{v}},{\bm{w}})$ lie within the annular rings $(r_{1},r_{2},r_{3})$ respectively. All these triangles have nearly the same shape and size. Although the triangles formed by $({\bm{u}},{\bm{v}},{\bm{w}})$ have different orientations in the plane perpendicular to $\hat{z}$, they all have nearly similar values of $(\mu_{1},\mu_{2},\mu_{3})$, or equivalently $\hat{\bm{p}}$ (Eq. 3), which decides the inclination with respect to $\hat{z}$. We use Eq. 2 and Eq. 3 to transform from $(k_{\perp r_{1}},k_{\perp r_{2}},k_{\perp r_{3}},k_{\parallel n_{1}},k_{\parallel n_{2}})$ to $(\hat{\textbf{p}},k_{1},\mu,t)$, whereby we have the binned bispectrum estimator $\hat{B}^{s}(\hat{\bm{p}},k_{1},\mu,t)$ that we use in Eq. 5 to calculate the multipole moments $\hat{\bar{B}}_{\ell}^{m}(k_{1},\mu,t)$.
We now briefly discuss the implementation of the bispectrum estimator. The definition of the annular rings is an important factor that affects the accuracy of the estimator, as well as its computational complexity. Here, we have chosen the ring width $(\delta k_{\perp})_{r}$, in grid units, to start from $1$ and double after every $10$ rings i.e., $(\delta k_{\perp})_{r}=2^{\lfloor(r-1)/10\rfloor}$, where $\lfloor$.$\rfloor$ . ⌋ denotes the floor function. As mentioned earlier, the largest ring is restricted to a maximum outer radius of $N_{g}/3$ by the periodic boundary condition imposed on the finite volume [35]. In this setting, $N_{r}$ has a logarithmic dependence on $N_{g}$ i.e. $N_{r}\sim\log(N_{g})$.
We have implemented the estimator in two steps. In Step I, we compute the function $D(r,n,\bm{x}_{\perp})$ (Eq. 9) and store it for each ring on all the planes. The computation time to compute a $D(r,n,\bm{x}_{\perp})$, which is a 2D FFT, scales as $\sim N_{g}^{2}\,\log(N_{g})$. Considering the $N_{r}$ rings and the $N_{g}$ planes involved in Step I, the computation time scales as $T_{\rm I}\sim N_{r}\times N_{g}\times N_{g}^{2}\,\log(N_{g})$. However, $N_{r}$ has a weak logarithmic dependence on $N_{g}$. Ignoring the logarithmic dependence, we have $T_{\rm I}\sim N_{g}^{3}$. We also compute and store $I(r,\bm{x}_{\perp})$ (Eq. 12) that does not vary with $n$, and it thus suffices to evaluate this for a single plane. The computation time for this scales as $N_{g}^{2}$, which is subdominant, and we ignore this contribution to $T_{\rm I}$.
In Step II, we consider all possible combinations of three rings and two planes. We implement Eq. 8 to compute $\hat{B}^{s}(\hat{\bm{p}},k_{1},\mu,t)$ for each distinct combination of rings and planes. The number of possible combinations scales as $N_{r}^{3}\times N_{g}^{2}$, and the sum $\sum_{\bm{x}_{\perp}}$ scales as $N_{g}^{2}$. Combining these factors, and ignoring the logarithmic dependence, we have $T_{\rm II}\sim N_{g}^{4}$.
The values of $(k_{1},\mu,t)$ at which we have the estimates of the binned bispectrum $\hat{B}^{s}(\hat{\bm{p}},k_{1},\mu,t)$ are not uniformly distributed. We have implemented a second level of binning to combine these to estimate the angular multipoles ( Eq. 5). We have divided the $k_{1}$ range into bins of equal logarithmic interval and also divided the range $0.5\leq\mu,t\leq 1$ into bins of equal linear interval. The exact number of bins that we have used here varies with $\ell$, and it has been adjusted to achieve a high signal-to-noise ratio without smearing out the shape and size dependence of the signal. Considering Eq. 5, we have implemented the second level of binning using
$\hat{\bar{B}}_{\ell}^{m}(\bar{k}_{1},\bar{\mu},\bar{t})=\sqrt{\frac{2\ell+1}{4\pi}}\frac{\sum[Y_{\ell}^{m}(\hat{\bm{p}})]^{*}\hat{B}^{s}(\hat{\bm{p}},k_{1},\mu,t)N_{\rm tri}}{\sum|Y_{\ell}^{m}(\hat{\bm{p}})|^{2}N_{\rm tri}}$ (13)
where $(\bar{k}_{1},\bar{\mu},\bar{t})$ in the l.h.s. refers to a particular interval in $(k_{1},\mu,t)$ respectively, and $\hat{\bar{B}}_{\ell}^{m}$ refers to the bin average estimate evaluated with respect to this interval. The sum in the r.h.s. refers to the estimates $\hat{B}^{s}(\hat{\bm{p}},k_{1},\mu,t)$ that occur in that particular interval. Note that $N_{\rm tri}$ (Eq. 11) depends on $(\hat{\bm{p}},k_{1},\mu,t)$, however, for the brevity of notation, we have not shown this explicitly.
In our implementation, we do not store the values of $\hat{B}^{s}(\hat{\bm{p}},k_{1},\mu,t)$ as the data volume is large. For each estimate of the binned bispectrum $\hat{B}^{s}(\hat{\bm{p}},k_{1},\mu,t)$, we identify the appropriate $(\bar{k}_{1},\bar{\mu},\bar{t})$ bin, and directly accumulate both the numerator and the denominator of Eq. 13 for the required values of $\ell$ and $m$.
Figure 2: Computation time for the bispectrum multipoles as a function of the number of grids $N_{g}$. The solid green line shows $T_{I}$ the time taken to perform all the FFT operations (Step I), and the solid orange line shows $T_{II}$ the time to compute the bispectrum multipoles (Step II) considering all possible closed triangles. The corresponding dashed lines show the expected scaling with $N_{g}$, as indicated in the figure legend. The solid black line shows the total number of the closed triangles $N_{\rm tri}$, and the dashed black line shows $N_{g}^{6}$. The computation has been performed on a single-core CPU.
We next discuss the computational complexity of the estimator. The computation time required for a direct estimate of the bispectrum multipoles by summing over ${\bm{k}}_{1}$ and ${\bm{k}}_{2}$ and identifying each closed triangle is expected to scale as $\sim N_{\rm tri}\propto N_{g}^{6}$. The solid black line in Fig. 2 shows $N_{\rm tri}$, the total number of triangles computed using Eq. 11, as a function of $N_{g}$. We see that this closely matches the expected $N_{\rm tri}\propto N_{g}^{6}$ behavior shown by the dashed black line. The solid green line shows $T_{I}$ the time taken to perform all 2D FFT operations in Eq. 9 and Eq. 12 and save them to disk. We see that this matches the expected $T_{I}\propto N_{g}^{3}$ behavior shown by the dashed green line. The solid orange line shows $T_{II}$ the time taken to compute Eq. 8 and Eq. 11 for all possible combinations of rings and planes, and subsequently use these to compute all the non-zero bispectrum multipole moments up to $\ell=6$ and $m=6$. We see that this matched the $T_{II}\propto N_{g}^{4}$ line shown by the dashed orange line. The slight deviation between the solid and dashed lines most probably arises from the $\log(N_{g})$ terms that have not been considered when considering the scaling behavior of the computation time.
We see (Fig. 2) that the total computation time is dominated by Step II for which $T_{II}\propto N_{g}^{4}$. The results for the computation time presented in this work have been performed on a single-core CPU. However, both Step I and Step II can be easily parallelized. Considering Step I, the 2D FFT operations in Eq. 9 and Eq. 12 for the different rings and planes can each be executed in parallel on a different thread. Taking into account Step II, Eq. 8 and Eq. 11 for all possible combinations of three rings and planes, each combination can be executed in parallel on a different thread. Ideally, we expect the computation time to be reduced approximately by the number of cores for a parallel implementation of the estimator. However, oftentimes the performance per core drops when using many concurrently, and the actual gain may be less than expected.
@@ III.1 Theoretical expressions
To validate the estimator, we generate a mildly non-Gaussian random field in redshift space $\delta^{s}(\bm{x})$ for which the bispectrum multipole moments are analytically derived. We start with generating a Gaussian random field in real space $\delta^{r}_{\rm G}(\bm{x})$ using a specified power spectrum $P(k)$ and introduce non-Gaussianity through a local quadratic bias,
@@
$B^{r}$ $(k_{1},k_{2},k_{3})$ (15)
$=2f_{\rm NG}[P(k_{1})P(k_{2})+P(k_{2})P(k_{3})+P(k_{3})P(k_{1})]~{}.$
The field is mapped to redshift space by applying the linear RSD,
@@
where $\Delta^{r}({\bm{k}})$ and $\Delta^{s}({\bm{k}})$ are the Fourier transform of the field in real and redshift space, respectively, and $\beta_{1}$ is the linear RSD parameter. The linear RSD theory predicts the redshift space bispectrum $B^{s}(\hat{\textbf{p}},k_{1},\mu,t)$ for this field to be,
$B^{s}$ $(\hat{\textbf{p}},k_{1},\mu,t)$ (17)
$=(1+\beta_{1}\mu_{1}^{2})(1+\beta_{1}\mu_{2}^{2})(1+\beta_{1}\mu_{3}^{2})B^{r}(k_{1},k_{2},k_{3})~{}.$
The analytical expressions for the multipole moments $B_{\ell}^{m}(k_{1},\mu,t)$ can be calculated using Eq. (4). These expressions are rather lengthy and, hence, not explicitly shown here. We refer the reader to Eqs. (24-29,31-33,A1-A7) of [25] for details. The analytical expressions are expected to hold true for sufficiently small values of $f_{\rm NG}$ such that $f_{\rm NG}\,\sigma_{\rm G}\ll 1$, where ${\sigma^{2}_{\rm G}=\langle[\delta^{r}_{\rm G}]^{2}\rangle}$ is the variance of ${\delta^{r}_{\rm G}}(\bm{x})$.
@@ III.2 Simulations
We simulated the field in a box of volume $V=L^{3}=(215~{}{\rm Mpc})^{3}$ divided in $N_{\rm g}=384$ grids in each direction, for which we have a grid spacing of $0.56~{}{\rm Mpc}$, and $k_{\rm min}=0.029~{}{\rm Mpc}^{-1}$. We use the input power spectrum $P(k)=k^{-2}$ to generate $\delta_{\rm G}^{r}$. The non-Gaussianity parameter is chosen to be $f_{\rm NG}=0.5$ for which $f_{\rm NG}\,\sigma_{\rm G}\approx 0.2$, and set the value of the linear RSD parameter $\beta_{1}=1$. We have generated $50$ independent realizations of the field and used these to estimate the mean $\bar{B}_{\ell}^{m}(\bar{k}_{1},\bar{\mu},\bar{t})=\langle\hat{\bar{B}}_{\ell}^{m}(\bar{k}_{1},\bar{\mu},\bar{t})\rangle$ and also the $1\sigma$ error bars.
@@ III.3 Binning and discreteness effects
In this work, we have validated the estimator by comparing the values of $\hat{\bar{B}}_{\ell}^{m}(\bar{k}_{1},\bar{\mu},\bar{t})$ estimated from the simulations with the analytical predictions for the same quantities. However, the analytical predictions $[\bar{B}_{\ell}^{m}]_{\rm c}$ presented in [25], which have been obtained by using Eq. (17) in Eq. (4), assume the continuum limit. In reality, we only have discrete ${\bm{k}}$ modes, and the integral in Eq. (4) is, in principle, replaced by the sum in Eq. (5). In the present work, we consider the binned bispectrum where the binning takes place in two steps. In the first step, we compute the binned bispectrum $\hat{B}^{s}(k_{\perp r_{1}},k_{\perp r_{2}},k_{\perp r_{3}},k_{\parallel n_{1}},k_{\parallel n_{2}})$ (Eq. 6). Here Eq. (8) directly yields the bispectrum averaged over a bin of triangles; these triangles all have nearly the same shape, size, and orientation with respect to $\hat{z}$. In the second step, we transform the variables $(k_{\perp r_{1}},k_{\perp r_{2}},k_{\perp r_{3}},k_{\parallel n_{1}},k_{\parallel n_{2}})\rightarrow(\hat{\textbf{p}},k_{1},\mu,t)$ and use Eq. (13) to calculate $\hat{\bar{B}}_{\ell}^{m}(\bar{k}_{1},\bar{\mu},\bar{t})$ where $(\bar{k}_{1},\bar{\mu},\bar{t})$ refers to a particular interval in $(k_{1},\mu,t)$ respectively, and $\hat{\bar{B}}_{\ell}^{m}$ refers to the average evaluated with respect to this interval. This average, which has $[Y_{\ell}^{m}(\hat{\bm{p}})]^{*}N_{\rm tri}$ as weights, combines triangles with nearly the same shape and size, but with different orientation with respect to $\hat{z}$.
To account for the effect of binning, we introduce the discrete analytical predictions $[\bar{B}_{\ell}^{m}]_{\rm d}(\bar{k}_{1},\bar{\mu},\bar{t})$ (as denoted by the subscript ${\rm d}$), which have been calculated by using Eq. (15) in Eq. (17). The resulting predictions have been binned using Eq. (13) in exactly the same way as for the estimator. Note that $[\bar{B}_{\ell}^{m}]_{\rm d}(\bar{k}_{1},\bar{\mu},\bar{t})$ only incorporates the binning in the second step and $[\bar{B}_{\ell}^{m}]_{\rm d}(\bar{k}_{1},\bar{\mu},\bar{t})$ does not incorporate the binning in the first step (Eq. 8). This would require us to perform the $\sim N_{g}^{6}$ computation that explicitly calculates the binned bispectrum (Eq. 6).
We validate the estimator by comparing $\bar{B}_{\ell}^{m}(k_{1},\mu,t)$ with the discrete analytical predictions $[\bar{B}_{\ell}^{m}]_{\rm d}(k_{1},\mu,t)$. Note that here and in the subsequent discussion, for brevity of notation, we have used $(k_{1},\mu,t)$ instead of $(\bar{k}_{1},\bar{\mu},\bar{t})$ as arguments for the final binned bispectrum.
Most theories which model the bispectrum typically provide analytical predictions for $[\bar{B}_{\ell}^{m}]_{\rm c}$. Therefore, it is often desirable to compare the estimated values $\bar{B}_{\ell}^{m}(k_{1},\mu,t)$ with the analytical predictions $[\bar{B}_{\ell}^{m}]_{\rm c}$ and not $[\bar{B}_{\ell}^{m}]_{\rm d}$. A few earlier works [32, 20] have addressed this issue by introducing ‘discreteness weights’ $w_{\ell}^{m}(k_{1},\mu,t)$
@@ III.4 Results
The $[\bar{B}_{\ell}^{m}]_{\rm c}$ values are predicted to have a scale-invariant power-law $k_{1}$ dependence, and it is thus adequate to show the results for just a single value of $k_{1}$. The number of triangles scales as $N_{\rm tri}\propto k_{1}^{6}$ [37], and we have chosen a large value $k_{1}=2.1~{}{\rm Mpc}^{-1}$ with a bin width $dk_{1}=0.2~{}{\rm Mpc}^{-1}$ so as to achieve a large value of $N_{\rm tri}$ and obtain a reliable (high SNR) estimate of $\bar{B}_{\ell}^{m}$. So as to accurately capture the shape dependence of $\bar{B}_{\ell}^{m}(k_{1},\mu,t)$, we have divided the $(\mu,t)$ plane into fine bins of width $(d\mu,dt)\approx(0.042,0.042)$ for evaluating the $\ell=0$ and $2$ multipoles. The values of $\bar{B}_{\ell}^{m}$ and the SNR are smaller for the higher multipoles $\ell>2$, and we have used coarser $(\mu,t)$ bins for these.
@@
Fig. 3 illustrates the $(\mu,t)$ bins that we have used to parameterize the triangle shape dependence of $\bar{B}_{\ell}^{m}(k_{1},\mu,t)$ for $\ell\leq 2$. It also shows $N_{\rm tri}$ the number of triangles corresponding to each bin. Considering the $\mu-t$ plane, the $\mu=1$ boundary corresponds to linear triangles where the two larger sides of the triangle are nearly aligned in the same direction. The $t=1$ and $12\mu t=$ 1 borders correspond to the L and S Isosceles triangles, respectively. The two largest sides are equal for L Isosceles triangles, whereas the two smaller sides are equal for S. We see that in all cases, the number of triangles is of the order of $10^{8}$, and it is largest for L Isosceles triangles. Further, $N_{\rm tri}$ decreases almost monotonically with the decline in $t$, and the number falls by a factor of $10$ as triangles are deformed to the stretched limit. Nonetheless, $N_{\rm tri}$ does not vary significantly with $\mu$.
@@
Figure 4: Results for $\ell=0$ and $2$ multipoles. It shows the estimates and their deviation from analytical prediction for all triangle shapes ($\mu,t$) having a fixed size $k_{1}=2.1~{}{\rm Mpc}^{-1}$. The uppermost row depicts the results for the monopole ($\ell=0,m=0$). The second, third, and fourth rows show $\ell=2$ and $m=0,1,2$ moments, respectively. The first column illustrates the mean estimates $\bar{B}_{\ell}^{m}$ from simulations. The second column shows $\Delta_{\ell}^{m}=|\Delta{B}_{\ell}^{m}|/{\sigma}_{B_{\ell}^{m}}$, $\Delta{B}_{\ell}^{m}$ is the deviation of the estimates from analytical predictions and ${\sigma}_{B_{\ell}^{m}}$ is the expected r.m.s. statistical fluctuation. The third column explicitly shows the estimated ${B}_{\ell}^{m}$ values (Red points with $5{\sigma}_{B_{\ell}^{m}}$ error bars) for linear, L, and S Isosceles triangles. The various lines show the corresponding analytical predictions. Results are plotted against $\mu$ for L and S Isosceles triangles and against $t$ for linear triangles.
@@
Fig. 4 shows the results for the $\ell=0$ and $2$ multipoles. Here, each row corresponds to a different combination of $\ell$ and $m$, as indicated in the figure. The panels in the left column show $\bar{B}_{\ell}^{m}$ estimated from the simulations. The analytical predictions $[\bar{B}_{\ell}^{m}]_{\rm d}$ are visually indistinguishable from the $\bar{B}_{\ell}^{m}$ shown here, and we have not explicitly shown $[\bar{B}_{\ell}^{m}]_{\rm d}$. The panels in the middle column show $\Delta_{\ell}^{m}=|\Delta{B}_{\ell}^{m}|/\sigma_{B_{\ell}^{m}}$ where $\Delta\bar{B}_{\ell}^{m}={\bar{B}}_{\ell}^{m}-[\bar{B}_{\ell}^{m}]_{\rm d}$ is the difference between the estimated value and the analytical predictions, and $\sigma_{B_{\ell}^{m}}$ (estimated from the $50$ realizations of the simulations) is the r.m.s. statistical fluctuations expected in $\Delta\bar{B}_{\ell}^{m}$. $\Delta_{\ell}^{m}$ provides an estimate of whether the deviations between the estimator and analytical predictions are consistent with the expected statistical fluctuations or not. The panels in the right column show a comparison of $\bar{B}_{\ell}^{m}$ and $[\bar{B}_{\ell}^{m}]_{\rm d}$ along the three boundaries of the $\mu-t$ plane.